A roof rises 2.75 feet over a horizontal distance of 10.69 feet. What is the slope of the roof to the nearest hundredth?
Found 2 solutions by atif.muhammad, tutorcecilia:
Answer by atif.muhammad(135)

This requires simple application of Pythagoras' theorem.
A roof rises 2.75 feet over a horizontal distance of 10.69 feet. What is the slope of the roof to the nearest hundredth?

Slope = Square root of (2.75^2 + 10.69^2) = 11.0380 = 11.04

Answer by tutorcecilia(2152)

Use the Pythagorian Theorem to solve:
a^2 + b^2 = c^2
a=10.69 [horizontal distance]
b=2.75 [rise of the roof]
c = unknown [slope of the roof, the hypothenuse]
.
(10.69)^2 + (2.75)^2 = c^2 [Plug-in the values]
114.2761 + 7.5625 = c^2 [Solve]
121.8386=c^2
sqrt121.8386 = sqrtc^2 [Take the square root of each side to find "c"]]
11.038 = 11.04 = c
.
Checkiing the answer:
(10.69)^2 + (2.75)^2 = (11.038)^2
121.830 = 121.838 Checks out.